# HeadLighter: Disentangling Illumination in Generative 3D Gaussian Heads via Lightstage Captures

### v1

@@ Abstract
Figure 1: Our method achieves high-fidelity disentanglement of intrinsic appearance and illumination in generative 3D Gaussian head models. It produces photorealistic intrinsic attributes including albedo, normals, and roughness, and supports relighting under diverse lighting conditions such as environment maps, spherical harmonics, and point lights. Left: intrinsic appearance attributes of two generated subjects. Right: relighting results of one subject under varying illumination.
1 Introduction
@@
Generative models for 3D human heads have seen remarkable progress, driven by their wide-ranging applications in digital avatars, virtual reality, and the entertainment industry. Among them, the most prominent approaches are 3D-aware Generative Adversarial Networks (GANs) built upon neural rendering techniques  [kerbl20233d, mildenhall2021nerf]. These models learn a rich prior of 3D head geometry and appearance directly from large-scale 2D in-the-wild images, without requiring any explicit 3D supervision. Consequently, they can synthesize highly realistic and view-consistent 3D head with rich identity and appearance diversity. In the pursuit of real-time performance and impressive visual quality, recent works integrate 3D Gaussian Splatting (3DGS) into the generative framework [gghead, barthel2025cgs] , replacing the computationally intensive volume rendering and achieving real-time rendering.
@@
However, a critical limitation persists: the illumination and the intrinsic head appearance (i.e., material or albedo) are deeply entangled, fundamentally preventing fine-grained control over the lighting conditions of the generated 3D heads. Existing approaches enable unsupervised decomposition of pretrained generative models using simplified reflectance models or strong assumptions. NFL  [nerffacelighting] involve style-mixing supervision with the white-light and face symmetry assumption, initializing from a on-the-shell light estimator, while GSHR [lv2025gsheadrelight] introduce a simplified light transport function for 3D gaussian and distill albedo and shading from NFL [nerffacelighting]. Others [deng2024lumigan, ranjan2023facelit] enforce physically plausible shading in neural volume rendering to split the diffuse and specular components. While these methods support explicit lighting control, their lack of real-world data limits their ability to handle complex lighting and achieve high physical plausibility.
@@
We address this by an intuitive and straight forward way: involving a multi-view face dataset captured under controlled illuminations with a LightStage to learn this disentanglement. However, directly utilizing this dataset is non-trivial: compared to large-scale in-the-wild face image datasets, our light stage dataset covers only a limited range of identities and lighting conditions, and exhibits a noticeable domain gap relative to natural imagery. To prevent overfitting on the light stage dataset and to preserve the generalization capability of the base generative model, we decompose pretrained generativte model into a dual-branch architecture: 1) a light-invariant base branch to generate geometry and albedo of 3D gaussian, and 2) a relight branch predicting attributes for physically-based rendering. And we further design a progressive training schema to ensure stable disentanglement. Firstly a pretrained generative model is shifted into light-invariant domain as the base branch to preserve rich generative prior of human heads. Guided by frozen base branch, we inject face appearance priors embedded in lightstage captures into the relight branch. Finally, two branches are trained jointly for better generalization.
@@
A further challenge lies in obtaining accurate surface normals, which are crucial for realistic shading but not native to 3D gaussian representations. While computing normals from extracted mesh or neighbor splats are feasible, such methods are computationally expensive in training iterations and can introduce significant noise. Observing that the base branch encodes gaussian geometry, we task the relight branch to directly predict plausible normals guided by base branch features. We further introduce a distillation strategy and smoothness terms that enforce consistent and plausible normal prediction, enabling photorealistic rendering of complex shading effects.
@@
We demonstrate the effectiveness of our pipeline through comprehensive comparisons with state-of-the-art methods. We evaluate generative diversity, rendering speed and multi-view spatial consistency of our disentangled generative models. We demonstrate the rendering results of 3D human heads generated by our method under different lighting inputs. To further validate the disentanglement capability, we quantitatively and qualitatively compare free-view relighting performance by extracting image pairs with varying poses and lighting conditions from our real-captured dataset. To foster future research, we will make our code and licensed dataset publicly available.
2 Related Works
2.1 3D-aware Generative Head Models
@@ 2 Related Works
The pursuit of high-fidelity, view-consistent, and controllable 3D human head synthesis has led to the rapid evolution of 3D-aware Generative Adversarial Networks. These models learn to generate 3D heads from large-scale 2D image sets, bypassing the need for large-scale 3D supervision. Early 3D GANs leverage various 3D representations [eg3d, chan2021pi, schwarz2020graf, next3d, tang20233dfaceshop] into the GAN framework. Among them, EG3D [eg3d] proposes a hybrid tri-planes, achieving photorealistic visual quality. But its reliance on volume rendering is computationally expensive, and super-resolution modules introduce view inconsistencies. To facilitate efficient, high-resolution generative head models, recent methods like [barthel2025cgs, yu25gaia, gghead] have turned to 3D Gaussian Splatting (3DGS) [kerbl20233d]. However, a fundamental limitation is that intrinsic face appearance properties and lighting are conflated into a view-dependent color, which makes direct and physically-plausible relighting impossible.
@@
To enable explicit lighting control, a line of research [pan2021shading, tan2022volux, nerffacelighting, deng2024lumigan, lv2025gsheadrelight, jiang2025nerffaceshop, ranjan2023facelit] has focused on disentangling illumination. NFL [nerffacelighting] distills pseudo albedo and shading from pretrained EG3D [eg3d] model by style-mixing, constrained by white-light, face symmetry assumption and an on-the shell light estimator. LumiGAN [deng2024lumigan] and FaceLit [ranjan2023facelit] employ an inverse rendering framework to separately model diffuse and specular components. GSHR [lv2025gsheadrelight] achieves highly efficient synthesis by designing a simplified radiance transfer representation for 3D Gaussians, but it relies on supervision from the pseudo-data generated by NFL [nerffacelighting]. Ultimately, these approaches lack access to ground-truth data, so they must resort to strong, simplifying assumptions that limit their ability to model the complex interaction between realistic materials and colored illumination. So we propose HeadLighter, a generative model that learns illumination disentanglement from lightstage captures and outputs high-fidelity relightable 3D gaussian heads.
2.2 Portrait Relighting
@@
Portrait relighting [yeh2022learning, hou2022face, pandey2021total, cai2024real, guo2025high, mei2023lightpainter, feng2021learning] aims to render a person’s face under novel lighting conditions. The core challenge lies in solving an ill-posed inverse rendering problem to separate intrinsic properties from the scene’s illumination. [debevec2000acquiring, rgca] capture One-Light-at-a-Time (OLAT) images with Lightstage and render under novel lighting, yielding high-fidelity relighting, but they require controlled capture setting and do not generalize to in-the-wild, sparse-view portraits. To enable relighting a single in-the-wild portrait, subsequent methods introduce various priors and assumptions. Early works [zhou2019deep, ponglertnapakorn2023difareli] rely on strong geometric priors, such as 3D Morphable Models [blanz2023morphable]. Others [wang2020single, chaturvedi2025synthlight, yeh2022learning] leverage synthetic datasets for training, but face a sim-to-real domain gap. Some works[cai2024real, Guo_2025_CVPR] support portrait video relighting which focus on consistency across frames.
@@
More recently, research has shifted towards leveraging the powerful generative priors. Initial attempts treat relighting as a style transfer task within the latent space of 2D GANs [abdal2021styleflow, deng2020disentangled, yi2024feditnet++, br2021photoapp, shoshan2021gan], but struggled to decouple illumination from other facial attributes like pose and expression. Recently, a growing body of work [lite2relight, holorelight, deng2024lumigan, pan2021shading, tan2022volux, nerffacelighting, lv2025gsheadrelight, rao20253dpr] imposes generative 3D face priors to jointly support relighting and novel-view synthesis. NFL [nerffacelighting] and GSHR [lv2025gsheadrelight] enable single portrait relighting through an inversion-then-relight pipeline using self-supervision. HoloRelight [holorelight] utilizes lightstage data and synthetic data for training which firstly delights the input and then applies a new lighting condition via an implicit shading representation. 3DPR [rao20253dpr] enables generating a basis of virtual OLAT images from a single portrait and then recombining them for relighting. Our work follows this direction. By integrating a physically-based lighting decomposition into a generative 3D Gaussian framework and training on lightstage data, HeadLighter combines the efficiency of explicit representations, the rich priors of generative models, and the physical accuracy of real-world measurements.
3 Methods
3.1 Preliminaries
@@
3DGS Head GAN leverages neural rendering techniques to learn priors of 3D facial shape and appearance from large-scale 2D image collections. To overcome the computational bottleneck of traditional volumetric rendering, recent works integrate the 3D GAN framework with fast rasterization of 3D Gaussian Splatting (3DGS) [kerbl20233d]. A prominent example is GGHead [gghead], which employs a StyleGAN2-based generator [karras2019style] to map a latent code $z\in\mathbb{R}^{512}$ to multi-channel 2D attribute maps $M_{\star}$ defined in the UV space of a template head mesh. The attributes for 3D Gaussian $\{G_{k}\}$, can be obtained by sampling the feature maps $M_{\star}$ at UV coordinate $x_{k}^{uv}$. These attributes are categorized into: (1) geometric properties, including 3D position $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$, rotation $\mathbf{q}_{k}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{k}\in\mathbb{R}^{3}$, and opacity $o_{k}\in\mathbb{R}$; and (2) view-dependent radiance defined with 1-degree Spherical Harmonics (SH) coefficients $a_{k}\in\mathbb{R}^{3\times(1+1)^{2}}$. The framework is further stabilized with several regularization terms during adversarial training to ensure high-fidelity results.
3.2 Data Acquisition
@@
We capture multi-view One-Light-A-Time using a Light Stage system equipped with 46 calibrated uniform distributed point lights and 28 synchronized cameras. Both the lights and cameras operate at 25 fps. We record under various lighting configurations, including a “Uniform” mode with all 46 lights turning on, a “Direction” mode where each light is activated together with its neighboring lights, and “Random” mode where 10 or 20 lights are randomly selected. To acquire robust 2D detection for dimly lit frames, we insert a fully illuminated frame every 10 frames, assume the subject remains stationary across these 11 frames, and propagate its estimation to the intervening frames. We capture three expressions for each subject: neutral, eyes closed, and smile. We provide more details about the dataset processing in 4.1. The captured images and processed data for participants with authorization will be publicly released.
3.3 Relightable 3DGS Head Generation
@@
We represent a relightable 3D head as a collection of 3D Gaussians $\{g_{k}\}$, where each Gaussian $g_{k}$ is associated with a set of geometric and appearance attributes and can be rendered into an image via splatting in a differentiable, physically-based manner, given arbitrary lighting conditions and viewpoints. All attributes of the 3D Gaussians are embedded into structured 2D UV maps of a template head mesh, from which each Gaussian $g_{k}$ samples its attributes at its assigned UV coordinate. We generate these maps using a dual-branch StyleGAN network that shares the common UV definition and similar network architecture:
@@
  • Base branch $G_{\text{base}}$ predicts the geometric attributes $\mathcal{A}_{\text{geo}}=(\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{q}_{k},o_{k})$ and albedo $\rho_{k}$.
@@
  • Relight branch $G_{\text{relit}}$ predicts the relighting-related attributes: surface normal $\mathbf{n}_{k}\in\mathbb{R}^{3}$, scalar roughness $\sigma_{k}\in\mathbb{R}$, and diffuse transport coefficients $\mathbf{T}_{k}\in\mathbb{R}^{3\times(s+1)^{2}}$, where $s$ is the degree of SH coefficients.
@@
Specifically, for the view-independent diffuse term $c_{\text{diffuse},k}$, we adopt Precomputed Radiance Transfer (PRT) [sloan2002precomputed] to model global illumination effects, including self-occlusion and subsurface scattering. The illumination from each light source is projected into SH coefficient vectors $\mathbf{L}_{i}\in\mathbb{R}^{(s+1)^{2}}$. The network predicts per-Gaussian diffuse transport coefficients $\mathbf{T}_{k}$, and the diffuse color is calculated by inner products in SH domain, modulated by the albedo $\rho_{k}$:
$c_{\text{diffuse},k}(\mathcal{L})=\rho_{k}\odot\sum_{i=1}^{M}\left(\mathbf{T}_{k}\cdot\mathbf{L}_{i}\right).$ (1)
The view-dependent specular term $c_{\text{specular},k}$ is modeled using a simplified Cook-Torrance microfacet model [cook1981reflectance] to create realistic highlights. For improved training stability, we retain only the core Normal Distribution Function (D) while omitting the Fresnel and Geometry terms. The total specular radiance is the sum of contributions from all $M$ light sources:
$c_{\text{specular},k}(\mathcal{L},\omega_{o})=\sum_{i=1}^{M}f_{\text{specular}}(\omega_{i},\omega_{o},\mathbf{n}_{k})(\mathbf{n}_{k}\cdot\omega_{i})\mathbf{I}_{i},$ (2)
@@
$f_{\text{specular}}(\omega_{i},\omega_{o},\mathbf{n}_{k})=\frac{D(\mathbf{h}_{i},\sigma_{k})}{4(\mathbf{n}_{k}\cdot\omega_{i})(\mathbf{n}_{k}\cdot\omega_{o})}.$ (3)
@@
Here, $\mathbf{h}_{i}=\frac{\omega_{i}+\omega_{o}}{\|\omega_{i}+\omega_{o}\|}$ is the halfway vector for the $i$-th light source. The $D$ term is modeled by the Trowbridge-Reitz GGX distribution parameterized by the per-Gaussian roughness $\sigma_{k}$:
$D(\mathbf{h}_{i})=\frac{\alpha_{k}^{2}}{\pi((\mathbf{n}_{k}\cdot\mathbf{h}_{i})^{2}(\alpha_{k}^{2}-1)+1)^{2}},\quad\alpha_{k}=\sigma_{k}^{2}.$ (4)
3.4 Disentanglement

### v2

@@ Abstract.
Generative Heads, Face Relighting
copyright: none
@@
Figure 1. Our method achieves high-fidelity disentanglement of intrinsic appearance and illumination in generative 3D Gaussian head models. It produces photorealistic intrinsic attributes including albedo/normals/roughness and utilizes precomputed radiance transfer (PRT) for diffuse global shading, and supports relighting under diverse lighting conditions such as environment maps, spherical harmonics. Left: intrinsic appearance attributes of two generated subjects (rendered under the SH lighting shown in the red boxes). Right: relighting results under varying environment maps(shown in the red boxes).
1. Introduction
@@
Generative models for 3D human heads have seen remarkable progress, driven by their wide applications in digital avatars and virtual reality. Among them, the most prominent approaches are 3D-aware Generative Adversarial Networks (GANs) built upon neural rendering techniques  (Kerbl et al., 2023; Mildenhall et al., 2021). These models learn a rich prior of 3D head geometry and appearance directly from large-scale 2D in-the-wild images, without requiring any explicit 3D supervision. Consequently, they can synthesize highly realistic and view-consistent 3D head with rich identity and appearance diversity. In the pursuit of real-time performance and impressive visual quality, recent works integrate 3D Gaussian Splatting (3DGS) into the generative framework (Kirschstein et al., 2024; Barthel et al., 2025) , replacing the computationally intensive volume rendering and achieving real-time rendering.
@@
However, a critical limitation persists: the illumination and the intrinsic head appearance (i.e., material or albedo) are deeply entangled, fundamentally preventing fine-grained control over the lighting conditions of the generated 3D heads. Existing approaches enable unsupervised decomposition of pretrained generative models using simplified reflectance models or strong assumptions. NFL  (Jiang et al., 2023) involves style-mixing supervision with the white-light and face symmetry assumption, initializing from an on-the-shell light estimator, while GSHR (Lv et al., 2025) introduces a simplified light transport function for 3D gaussian and distills albedo and shading from NFL (Jiang et al., 2023). Others (Deng et al., 2024; Ranjan et al., 2023) enforce physically plausible shading in neural volume rendering to split the diffuse and specular components. While these methods support explicit relighting, lack of real-world data supervision limits their ability to handle complex lighting and achieve highly physical plausibility.
@@
We address this by an intuitive and straightforward way: involving a multi-view face dataset captured under controlled illuminations with a LightStage to learn the disentanglement. However, directly utilizing such dataset is non-trivial: compared to large-scale in-the-wild face image datasets, our light stage dataset covers only a limited range of identities and lighting conditions, and exhibits a noticeable domain gap relative to natural imagery. To prevent overfitting on the lightstage dataset and preserve the generalization capability of the base generative model, we decompose pretrained generativte model into a dual-branch architecture: 1) a light-invariant base branch to generate geometry and albedo of 3D gaussian, and 2) a relight branch predicting attributes for physically-based rendering. To ensure stable disentanglement, we devise a progressive training strategy. First, we obtain base branch by fine-tuning the intermediate layers of a pretrained generator to remove illumination while preserving generation diversity. Next, with the base branch frozen, we inject head material priors from LightStage captures into the relight branch. Finally, we enhance the model’s generalization across diverse illumination environments by employing adversarial training with randomly sampled lighting conditions.
@@
A further challenge lies in obtaining accurate surface normals, which are crucial for realistic shading but not native to 3D gaussian representations. While computing normals from extracted mesh or neighbor splats are feasible, such methods are computationally expensive in training iterations and can introduce significant noise. Observing that the base branch encodes face geometry, we task the relight branch to directly predict normals guided by base branch features. We further introduce a distillation strategy and smoothness terms that enforce consistent and plausible normal prediction, enabling photorealistic rendering of complex shading effects.
@@
We demonstrate the effectiveness of our pipeline through comprehensive comparisons with state-of-the-art methods. We evaluate generative diversity, rendering speed and multi-view spatial consistency of our disentangled generative models. We demonstrate the rendering results of 3D human heads generated by our method under different lighting inputs. To further validate the disentanglement capability, we quantitatively and qualitatively compare free-view relighting results by extracting image pairs from our real-captured datasets. To foster future research, we will make our code and licensed dataset publicly available.
2. Related Works
2.1. 3D-aware Generative Head Models
@@ 2. Related Works
The pursuit of high-fidelity, view-consistent, and controllable 3D human head synthesis has led to the rapid evolution of 3D-aware Generative Adversarial Networks (GANs). These models learn to generate 3D heads from large-scale 2D image sets, bypassing the need for explicit 3D supervision. Early 3D GANs integrate various 3D representations (Chan et al., 2022, 2021; Schwarz et al., 2020; Sun et al., 2023; Tang et al., 2023) into the GAN framework. Among them, EG3D (Chan et al., 2022) utilizes a hybrid tri-plane representation, achieving photorealistic visual quality. However, its reliance on volume rendering is computationally expensive, and the super-resolution module can introduce view inconsistencies. To facilitate efficient, high-resolution generative head models, recent methods like (Barthel et al., 2025; Yu et al., 2025; Kirschstein et al., 2024) have turned to 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). However, a fundamental limitation persists: intrinsic face appearance properties and lighting are conflated into a view-dependent color, which renders direct and physically plausible relighting impossible.
@@
To enable explicit lighting control, a line of research (Pan et al., 2021; Tan et al., 2022; Jiang et al., 2023; Deng et al., 2024; Lv et al., 2025; Jiang et al., 2025; Ranjan et al., 2023) has focused on disentangling illumination. NFL (Jiang et al., 2023) distills pseudo albedo and shading from a pretrained EG3D (Chan et al., 2022) model via style-mixing, constrained by white-light and face symmetry assumptions, along with an on-the-shelf light estimator. LumiGAN (Deng et al., 2024) and FaceLit (Ranjan et al., 2023) employ an inverse rendering framework to separately model diffuse and specular components. GSHR (Lv et al., 2025) achieves highly efficient synthesis by designing a simplified radiance transfer representation for 3D Gaussians, but it relies on supervision from pseudo-data generated by NFL (Jiang et al., 2023). Consequently, these approaches lack access to ground-truth data and must resort to strong, simplifying assumptions that limit their ability to model the complex interaction between realistic materials and colored illumination. To address this, we propose HeadLighter, a high-fidelity relightable generative 3D Gaussian head model that learns illumination disentanglement from LightStage captures.
2.2. Portrait Relighting
@@
Portrait relighting (Yeh et al., 2022; Hou et al., 2022; Pandey et al., 2021; Cai et al., 2024; Guo et al., 2025b; Mei et al., 2023; Feng et al., 2021) aims to render a person’s face under novel lighting conditions. The core challenge lies in solving an ill-posed inverse rendering problem to separate intrinsic properties from the scene’s illumination. Existing works (Debevec et al., 2000; Saito et al., 2024) capture One-Light-at-a-Time (OLAT) images with LightStage to enable rendering under novel lighting, yielding high-fidelity results. However, they require controlled capture settings and do not generalize to in-the-wild, sparse-view portraits. To enable relighting of in-the-wild portraits, subsequent methods introduce various priors and assumptions. Early works (Zhou et al., 2019; Ponglertnapakorn et al., 2023; Fei et al., 2024) rely on low-rank priors like 3D Morphable Models (Blanz and Vetter, 2023). Others (Wang et al., 2020; Chaturvedi et al., 2025; Yeh et al., 2022) leverage synthetic datasets for training but face a sim-to-real domain gap. SwitchLight (Kim et al., 2024) learns diffusion prior from OLAT data and synthetic data to enable single-image portrait relighting. Similarly, URAvatar (Li et al., 2024) learns shape and appearance priors from large-scale OLAT data to support relightable reconstruction from smartphone videos. Some works (Cai et al., 2024; Guo et al., 2025a; Mei et al., 2025) address portrait video relighting, focusing on consistency across frames. Moreover, some methods also attempt relighting for both human body (Wang et al., 2025; Teufel et al., 2025) and general scenes (Zhang et al., 2025; Xing et al., 2025; Bharadwaj et al., 2025).
@@
Another line of research has shifted towards leveraging powerful generative priors. Initial attempts treated relighting as a style transfer task within the latent space of 2D GANs (Abdal et al., 2021; Deng et al., 2020; Yi et al., 2024; BR et al., 2021; Shoshan et al., 2021), but struggled to decouple illumination from other facial attributes like pose and expression. Recently, a growing body of work (Rao et al., 2024; Mei et al., 2024; Deng et al., 2024; Pan et al., 2021; Tan et al., 2022; Jiang et al., 2023; Lv et al., 2025; Rao et al., 2025) imposes generative 3D face priors to jointly support relighting and novel-view synthesis. HoloRelight (Mei et al., 2024) utilizes OLAT and synthetic data for training, first delighting the input and then applying new lighting conditions via an implicit shading representation. 3DPR (Rao et al., 2025) enables generating a basis of virtual OLAT images from a single portrait and then recombining them for relighting. Similar to NFL (Jiang et al., 2023) and GSHR (Lv et al., 2025), our method focuses on relightable 3D head generation but can also perform portrait relighting via an inversion-then-relight pipeline; however, under OLAT captures supervision, it achieves more physically accurate lighting disentanglement and thus better relighting quality.
3. Methods
3.1. Preliminaries
@@
3D Head GAN leverages neural rendering techniques to learn priors of 3D facial shape and appearance from large-scale 2D image collections. To overcome the computational bottleneck of traditional volumetric rendering, recent works integrate the 3D GAN framework with fast rasterization of 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023). A prominent example is GGHead (Kirschstein et al., 2024), which employs a StyleGAN2-based generator (Karras et al., 2019) to map a latent code $z\in\mathbb{R}^{512}$ to multi-channel 2D attribute maps $M_{\star}$ defined in the UV space of a template head mesh. The attributes for 3D Gaussian $\{G_{k}\}$, can be obtained by sampling the feature maps $M_{\star}$ at UV coordinate $x_{k}^{uv}$. These attributes are categorized into: (1) geometry properties, including position $\boldsymbol{\mu}_{k}\in\mathbb{R}^{3}$, rotation $\mathbf{q}_{k}\in\mathbb{R}^{4}$, scale $\mathbf{s}_{k}\in\mathbb{R}^{3}$, and opacity $o_{k}\in\mathbb{R}$; and (2) view-dependent radiance defined with 1-degree Spherical Harmonics (SH) $a_{k}\in\mathbb{R}^{3\times(1+1)^{2}}$. The framework is further stabilized with several regularization terms during adversarial training.
3.2. Data Acquisition
@@
We capture multi-view One-Light-A-Time(OLAT) images using LightStage system equipped with 46 calibrated uniform distributed point lights and 28 synchronized cameras. Both the lights and cameras operate at 25 fps. We record under various lighting configurations, including a “Uniform” mode with all 46 lights turning on, a “Direction” mode where each light is activated together with its neighboring lights, and “Random” mode where 10 or 20 lights are randomly selected. To acquire accurate 2D detection for each frames, we insert a fully illuminated frame every 10 frames, assume the subject remains stationary across these 11 frames, and propagate its estimation to the intervening frames. We capture three expressions for each subject: neutral, eyes closed, and smile. More details about the dataset processing can be found in 4.1.
3.3. Relightable 3DGS Head Generation
@@
We represent a relightable 3D head as a collection of 3D Gaussians $\{g_{k}\}$, where each Gaussian $g_{k}$ is associated with a set of geometric and appearance attributes and can be rendered into an image via splatting in a differentiable, physically-based manner, given arbitrary lighting conditions and viewpoints. All attributes of the 3D Gaussians are embedded into structured 2D UV maps of a template head mesh, from which each Gaussian $g_{k}$ samples its attributes at its assigned UV coordinate. We generate the geometry and material maps using a dual-branch StyleGAN generator. The two branches share a common UV parameterization but take distinct latent codes to disentangle geometry from lighting-specific attributes:
@@
  • Base branch $G_{\text{base}}:\mathbf{z}_{\text{base}}\to(\mathcal{A}_{\text{geo}},\boldsymbol{\rho})$. It maps a latent code $\mathbf{z}_{\text{base}}\in\mathbb{R}^{512}$ to the geometric attributes $\mathcal{A}_{\text{geo}}=\{\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{q}_{k},o_{k}\}$ and albedo $\boldsymbol{\rho}_{k}$.
@@
  • Relight branch $G_{\text{relit}}:\mathbf{z}_{\text{lit}}\to(\mathbf{n},\sigma,\mathbf{T})$. It takes a separate latent code $\mathbf{z}_{\text{lit}}\in\mathbb{R}^{128}$ to predict relighting-related attributes, including surface normals $\mathbf{n}_{k}\in\mathbb{R}^{3}$, scalar roughness $\sigma_{k}$, and diffuse light transport coefficients $\mathbf{T}_{k}\in\mathbb{R}^{3\times(s+1)^{2}}$, where $s$ denotes the SH degree.
@@
Specifically, for the view-independent diffuse term $c_{\text{diffuse},k}$, we adopt Precomputed Radiance Transfer (PRT) (Sloan et al., 2002) to model global illumination effects, including self-occlusion and subsurface scattering. The illumination from each light source is projected into SH coefficient vectors $\mathbf{L}_{i}\in\mathbb{R}^{(s+1)^{2}}$. The network predicts per-Gaussian diffuse transport coefficients $\mathbf{T}_{k}$, and the diffuse color is calculated by inner products in SH domain, modulated by the albedo $\rho_{k}$:
(1) $c_{\text{diffuse},k}(\mathcal{L})=\rho_{k}\odot\sum_{i=1}^{M}\left(\mathbf{T}_{k}\cdot\mathbf{L}_{i}\right).$
The view-dependent specular term $c_{\text{specular},k}$ is modeled using a simplified Cook-Torrance microfacet model (Cook and Torrance, 1981) to create realistic highlights. The total specular radiance is the sum of contributions from all $M$ light sources:
(2) $c_{\text{specular},k}(\mathcal{L},\omega_{o})=\sum_{i=1}^{M}f_{\text{specular}}(\omega_{i},\omega_{o},\mathbf{n}_{k})(\mathbf{n}_{k}\cdot\omega_{i})\mathbf{I}_{i},$
@@
(3) $f_{\text{specular}}(\omega_{i},\omega_{o},\mathbf{n}_{k})=\frac{D(\mathbf{h}_{i},\alpha_{k})\,F(\omega_{o},\mathbf{h}_{i})\,G(\omega_{i},\omega_{o},\alpha_{k})}{4\,(\mathbf{n}_{k}\cdot\omega_{i})\,(\mathbf{n}_{k}\cdot\omega_{o})},$
@@
Here, $\mathbf{h}_{i}=\frac{\omega_{i}+\omega_{o}}{\|\omega_{i}+\omega_{o}\|}$ is the halfway vector for the $i$-th light source. The $D$ term is modeled by the Trowbridge-Reitz GGX distribution (Trowbridge and Reitz, 1975) parameterized by the per-Gaussian roughness $\sigma_{k}$:
(4) $D(\mathbf{h}_{i})=\frac{\alpha_{k}^{2}}{\pi((\mathbf{n}_{k}\cdot\mathbf{h}_{i})^{2}(\alpha_{k}^{2}-1)+1)^{2}},\quad\alpha_{k}=\sigma_{k}^{2}.$
The Fresnel term $F$ is computed using Schlick’s approximation (Schlick, 1994)
(5) $F(\omega_{o},\mathbf{h}_{i})=F_{0}+(1-F_{0})(1-\max(\omega_{o}\cdot\mathbf{h}_{i},0))^{5},$
with a fixed dielectric Fresnel base reflectance $F_{0}=0.04$ for the entire head, consistent with skin and hair materials. The geometry term $G$ follows the Smith masking-shadowing model for GGX as described in (Walter et al., 2007).
3.4. Disentanglement